40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4792     Entry time: Mon Jun 6 23:56:16 2011
Author: rana, valera 
Type: Configuration 
Category: SUS 
Subject: ETMX/ETMY OSEM whitening 

 We measured the OSEM PD whitening transfer function of the ETMX OSEM UL whitening stage (D000210) by comparing the input signal to the whitening amplifier (single pin LEMO monitor) to the output signal - both were piped into the DAQ. The transfer function was close to constant 0 dB/180 deg independent of the whitening switch selection (FM1 filter engaged/disengaged)  up to ~20 Hz where we run out of coherence. All other ETMX and ETMY spectra at the input of the digital whitening compensation don't change when the whitening is switched on/off so by induction we conclude that all the ETMX/ETMY OSEM PD hardware whitening filters are not on.

ELOG V3.1.3-